1. A disk with a diameter of 18 cm is spun at a constant rate. The disk rotates 9 times every 1.3 seconds
a) What are the frequency and period of rotation?
b) What is the angular velocity in radians/sec?
c) What is the tangential velocity of a point on the edge of the disk?
1
Expert's answer
2019-10-31T11:18:21-0400

a) Frequency of rotation, f


"f=\frac{9 cycles}{1.3 seconds} = 6.9\frac{cycles}{second} = 6.9 Hz"

Period of rotation, T



"T =\frac{1}{f} = \frac {1}{6.9 Hz} = 0.14 s"

b) Angular velocity, ω



"ω = 2πf = 2\times 3.14\times 6.9 Hz = 43 \frac {radians}{sec}"

c) tangential velocity of a point on the edge of the disk, ν



"ν=ωr = ω\times \frac {d}{2} = 43 \frac {radians}{sec}\times\frac{18 cm}{2} = 387 \frac{cm}{s}"